OpenStudy (anonymous):

A long, straight wire of radius R carries a current I that is uniformly distributed over the cross-sectional area of the wire. Find the magnetic field both outside and inside the wire.

OpenStudy (anonymous):

You may need three equations to calculate the magnetic field in and around the conductor with radius ‘R’ and current ‘I’. Consider ‘r’ is the distance from the center of the wire where we have to measure the magnetic field. 1. if r<R then H=I*r/2*pi*R^2 2. if r=R then H=I/2*pi*R 3. if r>R then H=I/2*pi*r H=Magnetic field